# Leave-One-Out Analysis for Nonconvex Robust Matrix Completion with General Thresholding Functions

### v1

@@ 5 Induction Steps
$\leq$ $K\left\|\mathcal{P}_{\Omega^{\prime}}\left(L^{t,m}-L^{t}\right)\right\|_{\mathrm{F}}$
$\leq$ $K\left\|\mathcal{P}_{\Omega^{\prime}}\left(D_{U}^{t,m,0}\Sigma^{t,m}\left(V^{t,m}\right)^{T}\right)\right\|_{\mathrm{F}}+K\left\|\mathcal{P}_{\Omega^{\prime}}\left(U^{t}S^{t,m,0}\left(V^{t,m}\right)^{T}\right)\right\|_{\mathrm{F}}$
$+K\left\|\mathcal{P}_{\Omega^{\prime}}\left(U^{t}\Sigma^{t}\left(D_{V}^{t,0,m}\right)^{T}\right)\right\|_{\mathrm{F}}$
$\leq$ $K\sqrt{40\alpha^{\prime}p\kappa}\cdot\left(\left\|D_{U}^{t,m,0}\Sigma^{t,m}\left(V^{t,m}\right)^{T}\right\|_{\mathrm{F}}+\left\|U^{t}S^{t,m,0}\left(V^{t,m}\right)^{T}\right\|_{\mathrm{F}}+\left\|U^{t}\Sigma^{t}\left(D_{V}^{t,0,m}\right)^{T}\right\|_{\mathrm{F}}\right)$
$\leq$ $K\sqrt{40\alpha^{\prime}p\kappa}\cdot\left(2\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\left\|D^{t,m,0}\right\|_{\mathrm{F}}+\left\|S^{t,m,0}\right\|_{\mathrm{F}}\right),$
where the first inequality follows from property P.2 of the thresholding function (applied to each entry in $\Omega_{S^{\star}}^{(-m)}$), and the fourth inequality follows from Lemma 13 again, due to the fact that $D_{U}^{t,m,0}\Sigma^{t,m}\left(V^{t,m}\right)^{T}$ and $U^{t}S^{t,m,0}\left(V^{t,m}\right)^{T}$ are in the tangent space at
$U^{t,m}\left(G^{t,m}\right)^{T}\cdot\Sigma^{t,m}\cdot\left(V^{t,m}\left(G^{t,m}\right)^{T}\right)^{T},$
and $U^{t}\Sigma^{t}\left(D_{V}^{t,0,m}\right)^{T}$ is in the tangent space at
$U^{t}\left(G^{t}\right)^{T}\cdot\Sigma^{t}\cdot\left(V^{t}\left(G^{t}\right)^{T}\right)^{T}.$
Note that their closeness to the ground truth factors can be similarly verified as in (18).
@@
This bound is established in (20) for the induction steps, and for the base case, the bound (10) can also be written in this form. Applying Lemma 9 with $A:=\widehat{L^{\star}}+\widehat{E^{t-1,m}}$ and $\widetilde{A}:=\widehat{L^{\star}}+\widehat{E^{t-1}}$ yields
@@
$B_{1}\leq$ $p^{-1}\cdot\sqrt{40\alpha^{\prime}p\kappa r}\cdot\left\|\mathcal{P}_{\Omega_{S^{\star}}}^{(-m)}\left(S^{t,m}-S^{t}\right)\right\|_{\mathrm{F}}$
$\leq$ $40K\cdot\alpha^{\prime}\kappa\sqrt{r}\cdot\left(\left(2+\frac{2}{C_{0}}\right)\sigma_{1}^{\star}\left\|D^{t,m,0}\right\|_{\mathrm{F}}+\left\|S^{t,m,0}\right\|_{\mathrm{F}}\right)\leq\frac{C_{10}}{5C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}$
@@
$\alpha^{\prime}\leq\frac{1}{\kappa^{2}r^{1/2}}\cdot\frac{\gamma}{5000K}.$
@@
$B_{2}\leq$ $p^{-1}\cdot(2\alpha pn)\cdot C_{\text{thresh}}\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}\cdot\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)\leq\frac{1}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa^{2}\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$
@@
For $\tilde{B}_{1}$, one has
$\tilde{B}_{1}\leq$ $\left\|\mathcal{H}_{\Omega}\left(D_{U}^{t,m,0}\Sigma^{t,m}\left(V^{t,m}\right)^{T}\right)\Delta_{V}^{t+1,m}\right\|_{\mathrm{F}}+\left\|\mathcal{H}_{\Omega}\left(D_{U}^{t,m,0}\Sigma^{t,m}\left(V^{t,m}\right)^{T}\right)V^{\star}\right\|_{\mathrm{F}}$
$\leq$ $\underbrace{\left\|\mathcal{H}_{\Omega}\left(D_{U}^{t,m,0}\Sigma^{t,m}\left(V^{t,m}\right)^{T}\right)\Delta_{V}^{t+1,m}\right\|_{\mathrm{F}}}_{\tilde{B}_{1a}}+\underbrace{\left\|\mathcal{H}_{\Omega}\left(D_{U}^{t,m,0}\Sigma^{t,m}\left(\Delta_{V}^{t,m}\right)^{T}\right)V^{\star}\right\|_{\mathrm{F}}}_{\tilde{B}_{1b}}$
$+\underbrace{\left\|\mathcal{H}_{\Omega}\left(D_{U}^{t,m,0}\Sigma^{t,m}\left(G^{t,m}\right)^{T}\left(V^{\star}\right)^{T}\right)V^{\star}\right\|_{\mathrm{F}}}_{\tilde{B}_{1c}}.$
The application of Lemma 11 yields that
$\tilde{B}_{1a}\leq$ $\left\|D_{U}^{t,m,0}\Sigma^{t,m}\right\|_{\mathrm{F}}\cdot\max_{1\leq j\leq n}\left\|R^{(j)}\right\|_{2},$
@@
$R^{(j)}:=\sum_{k=1}^{n}\left(1-\delta_{jk}/p\right)\left(V^{t,m}_{k,:}\right)^{T}\left(\Delta_{V}^{t+1,m}\right)_{k,:}$
On the event that $\sum_{k=1}^{n}\delta_{jk}\leq 2pn,$ we have the direct calculation
$\left\|R^{(j)}\right\|_{2}\leq$ $3n\left\|V^{t,m}\right\|_{2,\infty}\left\|\Delta_{V}^{t+1,m}\right\|_{2,\infty}$
$\leq$ $3n\cdot\left(1+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\cdot\frac{5C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}\leq\frac{\gamma}{135\kappa}$
provided that $C_{0}\geq 2700C_{10}$. Therefore,
$\tilde{B}_{1a}\leq\frac{8C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\cdot\frac{\gamma}{135\kappa}\leq\frac{C_{10}}{15C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$
The term $\tilde{B}_{1b}$ can be bounded in the same way, and a same bound can be obtained as $\tilde{B}_{1a}$. For $\tilde{B}_{1c}$, one has
$\tilde{B}_{1c}\leq$ $\left\|D_{U}^{t,m,0}\Sigma^{t,m}\right\|_{\mathrm{F}}\cdot\max_{1\leq j\leq n}\left\|R^{(j)}\right\|_{2},$
where, with a slight abuse of notation,
$R^{(j)}:=\sum_{k=1}^{n}\left(1-\delta_{jk}/p\right)\left(V^{\star}_{k,:}\right)^{T}V^{\star}_{k,:}.$
Similar to the bound of (16), applying Lemma 10,
$\left\|R^{(j)}\right\|_{2}\leq$ $C_{10}\cdot\left(\sqrt{\frac{\log n}{p}}\left\|V^{\star}\right\|_{2,\infty}+\frac{\log n}{p}\left\|V^{\star}\right\|_{2,\infty}^{2}\right)\leq 2C_{10}\sqrt{\frac{\mu r\log n}{np}}\leq\frac{\gamma}{135\kappa}$
provided that
$p\geq\frac{72900C_{10}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{2}\mu r\log n}{n}.$
Therefore, $\tilde{B}_{1c}$ also has the same bound as $\tilde{B}_{1a}$. Putting them together, one has
$\tilde{B}_{1}\leq\frac{C_{10}}{5C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$
@@
$\tilde{B}_{3}\leq$ $\left\|\mathcal{H}_{\Omega}\left(\bm{1}\bm{1}^{T}\right)\right\|_{2}\left\|U^{t}\Sigma^{t}\right\|_{2,\infty}\left\|V^{t+1,m}\right\|_{2,\infty}\left\|D_{V}^{t,0,m}\right\|_{\mathrm{F}}$
$\leq$ $c_{5}\sqrt{\frac{n\log n}{p}}\cdot\left(1+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\cdot\left(1+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\cdot\frac{8C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}$
$\leq$ $\frac{1}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1},$
@@
$p\geq\frac{100c_{5}^{2}C_{10}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{2}\mu^{2}r^{2}\log n}{n}.$
@@
$B_{4}=\left\|e_{m}^{T}\mathcal{H}_{\Omega}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}\right\|_{2}\leq\frac{C_{10}}{5C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1},$
@@
$p\geq\frac{14400C_{10}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{2}\mu^{2}r^{2}\log n}{n}.$
@@
$\leq$ $\left\|\mathcal{H}_{\Omega}\left(L^{t,m}-L^{\star}\right)\right\|_{2}\left\|V^{t+1,m}\right\|_{2,\infty}\leq\frac{1}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa^{2}\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1},$
@@
$\left\|\left(E^{t}-E^{t,m}\right)V^{t+1,m}\right\|_{\mathrm{F}}\leq$ $\left(4\times\frac{C_{10}}{5C_{0}}+2\times\frac{1}{C_{0}}\right)\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}\leq\frac{C_{10}}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1},$
@@
$\left\|W^{t,m}F^{t+1,m}\right\|_{\mathrm{F}}\leq\frac{2C_{10}}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$ (20)
@@
$\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}\leq 2\cdot\max_{m}\left\|D^{t+1,0,m}\right\|_{\mathrm{F}}\leq\frac{8C_{10}}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$
@@
$C_{10}\geq 20,~{}C_{0}\geq\max\left\{2700C_{10},\frac{6C_{10}}{c_{13}}\right\},$ (21)
where $c_{13}>0$ is the small universal constant from Lemma 13, and $p$ and $\alpha$ satisfy
$p\geq\frac{\max\left\{16c_{4}^{2}C_{0}^{2}\cdot\kappa^{6}\mu^{3}r^{3},~{}1152c_{5}^{2}C_{10}^{2}\cdot\kappa^{4}\mu^{2}r^{3},~{}14400C_{10}^{2}\cdot\kappa^{2}\mu^{2}r^{2},~{}72900C_{10}^{2}\cdot\kappa^{2}\mu r\right\}}{\gamma^{2}}\cdot\frac{\log n}{n},$
$\alpha\leq\min\left\{\frac{1}{4C_{0}}\frac{1}{\kappa^{3}\mu^{2}r^{2}}\cdot\frac{\gamma}{C_{\text{thresh}}},~{}\frac{1}{\kappa^{2}r^{1/2}}\cdot\frac{\gamma}{5000K}\right\}.$

### v2

@@ 5 Induction Steps
$\leq$ $K\left\|\mathcal{P}_{\Omega_{S^{\star}}}\left(L^{t,m}-L^{t}\right)\right\|_{\mathrm{F}}$
$\leq$ $K\left\|\mathcal{P}_{\Omega_{S^{\star}}}\left(D_{U}^{t,m,0}\Sigma^{t,m}\left(V^{t,m}\right)^{T}\right)\right\|_{\mathrm{F}}+K\left\|\mathcal{P}_{\Omega_{S^{\star}}}\left(U^{t}S^{t,m,0}\left(V^{t,m}\right)^{T}\right)\right\|_{\mathrm{F}}$
$+K\left\|\mathcal{P}_{\Omega_{S^{\star}}}\left(U^{t}\Sigma^{t}\left(D_{V}^{t,0,m}\right)^{T}\right)\right\|_{\mathrm{F}}$
$\leq$ $K\sqrt{2\alpha pn}\cdot\left(\left\|D_{U}^{t,m,0}\Sigma^{t,m}\right\|_{\mathrm{F}}\left\|V^{t,m}\right\|_{2,\infty}+\left\|U^{t}S^{t,m,0}\right\|_{\mathrm{F}}\left\|V^{t,m}\right\|_{2,\infty}+\left\|U^{t}\right\|_{2,\infty}\left\|D_{V}^{t,0,m}\Sigma^{t}\right\|_{\mathrm{F}}\right)$
$\leq$ $K\sqrt{2\alpha pn}\cdot\left(2\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\left\|D^{t,m,0}\right\|_{\mathrm{F}}+\left\|S^{t,m,0}\right\|_{\mathrm{F}}\right)\cdot\left(1+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}},$
where the first inequality follows from property P.2 of the thresholding function (applied to each entry in $\Omega_{S^{\star}}^{(-m)}$), and the fourth inequality follows from Lemma 13 again.
@@
For the base case, such a bound can be established by substituting the bound of $\left\|\Delta^{1,\infty}\right\|_{2,\infty}$ into (10). For the induction steps, such a bound can be established by substituting the bound of $\left\|\Delta^{t,\infty}\right\|_{2,\infty}$ and $\left\|D^{t,\infty}\right\|_{\mathrm{F}}$ into (20) (the numbering is changed from $t+1$ to $t$). Applying Lemma 9 with $A:=\widehat{L^{\star}}+\widehat{E^{t-1,m}}$ and $\widetilde{A}:=\widehat{L^{\star}}+\widehat{E^{t-1}}$ yields
@@
$B_{1}\leq$ $p^{-1}\cdot\left\|\mathcal{P}_{\Omega_{S^{\star}}}^{(-m)}\left(S^{t,m}-S^{t}\right)\right\|_{\mathrm{F}}\cdot\sqrt{2\alpha pnr}\left\|V^{t+1,m}\right\|_{2,\infty}$
$\leq$ $4K\cdot\left(\left(2+\frac{2}{C_{0}}\right)\sigma_{1}^{\star}\left\|D^{t,m,0}\right\|_{\mathrm{F}}+\left\|S^{t,m,0}\right\|_{\mathrm{F}}\right)\cdot\alpha r\sqrt{\mu n}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $100K\cdot\frac{C_{10}}{C_{0}}\left(\frac{\sigma_{1}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\alpha r\sqrt{\mu n}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $\frac{C_{10}}{20C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\right)\gamma^{t+1}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
@@
$\alpha\leq\frac{1}{\kappa\mu r^{3/2}}\cdot\frac{\gamma}{2000K}.$
@@
$B_{2}\leq$ $p^{-1}\cdot(2\alpha pn)\cdot C_{\text{thresh}}\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}\cdot\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $\frac{1}{2C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\right)\gamma^{t+1}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right).$
@@
For $\tilde{B}_{1}$, the application of Lemma 11 yields that
$\tilde{B}_{1}\leq$ $\left\|D_{U}^{t,m,0}\Sigma^{t,m}\right\|_{\mathrm{F}}\cdot\max_{1\leq j\leq n}\left\|R^{(j)}\right\|_{2},$
@@
$R^{(j)}:=\sum_{k=1}^{n}\left(1-\delta_{jk}/p\right)\left(V^{t,m}_{k,:}\right)^{T}V^{t+1,m}_{k,:}.$
For each $j\in\{1,\cdots,n\}$, denoting $H^{(j)}:=\operatorname{diag}\left(1-\delta_{j1}/p,\cdots,1-\delta_{jn}/p\right)\in\mathbb{R}^{n\times n}$,
$\left\|R^{(j)}\right\|_{2}=$ $\left\|\left(V^{t,m}\right)^{T}H^{(j)}V^{t+1,m}\right\|_{2}$
$=$ $\left\|\left(D_{V}^{t,m,j}+V^{t,j}G^{t,m,j}\right)^{T}H^{(j)}\left(D_{V}^{t+1,m,j}+V^{t+1,j}G^{t+1,m,j}\right)\right\|_{2}$
$\leq$ $\left\|\left(D_{V}^{t,m,j}\right)^{T}H^{(j)}D_{V}^{t+1,m,j}\right\|_{2}+\left\|\left(V^{t,j}\right)^{T}H^{(j)}D_{V}^{t+1,m,j}\right\|_{2}$
$+\left\|\left(D_{V}^{t,m,j}\right)^{T}H^{(j)}V^{t+1,j}\right\|_{2}+\left\|\left(V^{t,j}\right)^{T}H^{(j)}V^{t+1,j}\right\|_{2}.$
Conditioning on the event that $\sum_{k=1}^{n}\delta_{jk}\leq 2pn,$
$\left\|\left(D_{V}^{t,m,j}\right)^{T}H^{(j)}D_{V}^{t+1,m,j}\right\|_{2}=$ $\left\|\sum_{k=1}^{n}\left(1-\delta_{jk}/p\right)\left(D_{V}^{t,m,j}\right)_{k,:}^{T}\left(D_{V}^{t+1,m,j}\right)_{k,:}\right\|_{2}$
$\leq$ $3n\left\|D_{V}^{t,\infty}\right\|_{2,\infty}\left\|D_{V}^{t+1,\infty}\right\|_{2,\infty}$
$\leq$ $3n\cdot\frac{8C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}\leq\sqrt{\mu rn}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}},$
$\left\|\left(V^{t,j}\right)^{T}H^{(j)}D_{V}^{t+1,m,j}\right\|_{2}\leq$ $3n\left\|V^{t,j}\right\|_{2,\infty}\left\|D_{V}^{t+1,\infty}\right\|_{2,\infty}$
$\leq$ $3n\cdot\left(1+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}\leq 4\sqrt{\mu rn}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}},$
$\left\|\left(D_{V}^{t,m,j}\right)^{T}H^{(j)}V^{t+1,j}\right\|_{2}\leq$ $3n\left\|D_{V}^{t,\infty}\right\|_{2,\infty}\left\|V^{t+1,j}\right\|_{2,\infty}$
$\leq$ $3n\cdot\frac{8C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $\frac{\gamma}{360\kappa}\sqrt{\frac{n}{\mu r}}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
provided that $C_{0}\geq 8640C_{10}$.
Since $V^{t,j}$ and $V^{t+1,j}$ are independent with respect to the random variables on the $j$-th row,
$\left\|\left(V^{t,j}\right)^{T}H^{(j)}V^{t+1,j}\right\|_{2}=$ $\left\|\sum_{k=1}^{n}\left(1-\delta_{jk}/p\right)\left(V^{t,j}_{k,:}\right)^{T}V^{t+1,j}_{k,:}\right\|_{2}$
$\leq$ $\frac{\gamma}{360\kappa}\sqrt{\frac{n}{\mu r}}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
where the last bound follows from the same argument for (16), provided that
$p\geq\frac{640000C_{10}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{2}\mu r\log n}{n}.$
Therefore,
$\tilde{B}_{1}\leq$ $\left\|D_{U}^{t,m,0}\Sigma^{t,m}\right\|_{\mathrm{F}}\left[\frac{\gamma}{180\kappa}\sqrt{\frac{n}{\mu r}}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)+5\sqrt{\mu rn}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}\right]$
$\leq$ $\frac{8C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\left[\frac{\gamma}{180\kappa}\sqrt{\frac{n}{\mu r}}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)+5\sqrt{\mu rn}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}\right]$
$\leq$ $\frac{C_{10}}{20C_{0}}\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{C_{10}}{20C_{0}}\frac{\sigma_{r}^{\star}}{\kappa}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}+\frac{\sigma_{r}^{\star}}{32}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}.$
@@
$\tilde{B}_{3}\leq$ $\left\|\mathcal{H}_{\Omega}\left(\bm{1}\bm{1}^{T}\right)\right\|_{2}\left\|U^{t}\Sigma^{t}\right\|_{2,\infty}\left\|D_{V}^{t,0,m}\right\|_{\mathrm{F}}\left\|V^{t+1,m}\right\|_{2,\infty}$
$\leq$ $c_{5}\sqrt{\frac{n\log n}{p}}\cdot\left(1+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\cdot\frac{8C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $\frac{1}{2C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\right)\gamma^{t+1}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$p\geq\frac{400c_{5}^{2}C_{10}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{2}\mu^{2}r^{2}\log n}{n}.$
@@
$B_{4}=\left\|e_{m}^{T}\mathcal{H}_{\Omega}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}\right\|_{2}\leq\frac{C_{10}}{20C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\right)\gamma^{t+1}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$p\geq\frac{230400C_{10}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{4}\mu^{2}r^{2}\log n}{n}.$
@@
$\leq$ $\left\|\mathcal{H}_{\Omega}\left(L^{t,m}-L^{\star}\right)\right\|_{2}\left\|V^{t+1,m}\right\|_{2,\infty}\leq\frac{1}{2C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\right)\gamma^{t+1}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$\left\|\left(E^{t}-E^{t,m}\right)V^{t+1,m}\right\|_{\mathrm{F}}\leq$ $\left(4\times\frac{C_{10}}{20C_{0}}+2\times\frac{1}{2C_{0}}\right)\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}$
$+\left(4\times\frac{C_{10}}{20C_{0}}+2\times\frac{1}{2C_{0}}\right)\frac{\sigma_{r}^{\star}}{\kappa}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}+\frac{\sigma_{r}^{\star}}{16}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}},$
@@
$\left\|W^{t,m}F^{t+1,m}\right\|_{\mathrm{F}}\leq\frac{C_{10}}{2C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{\sigma_{r}^{\star}}{10\kappa}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}+\frac{\sigma_{r}^{\star}}{8}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}.$ (20)
@@
$\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}\leq 2\cdot\max_{m}\left\|D^{t+1,0,m}\right\|_{\mathrm{F}}\leq\frac{2C_{10}}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{2}{5\kappa}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}+\frac{1}{2}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}},$
and as a result,
$\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}\leq\frac{4C_{10}}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{4}{5\kappa}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}.$ (21)
Combining (21) with (17),
$\left(1-\frac{4+3C_{10}}{C_{0}}\right)\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\leq\frac{21+3C_{10}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{1}{C_{0}}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty},$
$\left(1-\frac{5+3C_{10}}{C_{0}}\right)\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\leq\frac{21+3C_{10}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$
When $C_{0}\geq 5(5+3C_{10})$ and $C_{10}\geq 21$,
$\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\leq\frac{5C_{10}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1},$
substituting this bound into (21) and we get
$\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}\leq\frac{8C_{10}}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$
@@
$C_{10}\geq 21,~{}C_{0}\geq 8640C_{10},$ (22)
and $p$ and $\alpha$ satisfy
$p\geq\frac{\max\left\{16c_{4}^{2}C_{0}^{2}\cdot\kappa^{4}\mu^{3}r^{3},~{}1152c_{5}^{2}C_{10}^{2}\cdot\kappa^{4}\mu^{2}r^{3},~{}230400C_{10}^{2}\cdot\kappa^{4}\mu^{2}r^{2},~{}640000C_{10}^{2}\cdot\kappa^{2}\mu r\right\}}{\gamma^{2}}\cdot\frac{\log n}{n},$
$\alpha\leq\min\left\{\frac{1}{4C_{0}}\frac{1}{\kappa^{2}\mu^{2}r^{2}}\cdot\frac{\gamma}{C_{\text{thresh}}},~{}\frac{1}{\kappa\mu r^{3/2}}\cdot\frac{\gamma}{2000K}\right\}.$